### v1

@@ 3.3 The chiral magnetic effect
Using a plane-parallel model (Lander and Gourgouliatos, 2019) investigated the features of such plastic flow under the assumption of Stokes flow, where a viscous term balances magnetic and elastic stresses. They study the crustal response under Ohmic and Hall evolution and find that there can be significant plastic-like motions in the external layers of the star. Similar arguments have also been proposed to account for the deposition of heat by the visco-plastic flow and the propagation of thermo-plastic waves (Beloborodov and Levin, 2014). In a later work, Gourgouliatos and Lander (2021) conducted global axisymmetric simulations to investigate various failure mechanisms impacting a specific crustal region. They find that flow does not merely dampen the Hall effect, even when plastic viscosity is low; instead, it drives complex evolutionary patterns, sometimes amplifying the Hall effect’s influence. They concluded that plastic flow influences both the characteristics of magnetar bursts and their spin-down behavior.
However, as pointed out in recent work (Kojima, 2024; Bransgrove et al, 2025), in magneto-elastic equilibrium, it is the divergence of stress (i.e., the force) that is balanced, not the stress itself. Kojima’s methodology for calculating the elastic tensor is needed to ensure the model accuracy. In numerical simulations of magnetic field evolution, $\sigma_{ij}$ must be tracked by solving additional differential equations, at the cost of a significant increase in computation time compared to previous simple algebraic relations. Bransgrove et al (2025) indicate that, in configurations close to magneto-elastic equilibrium, the elastic stress is typically several orders of magnitude smaller than the local Maxwell stress. Consequently, criteria based only on stress balance may significantly overestimate the frequency of crustal failures, and previous conclusions should be approached with caution. Nevertheless, the results in Bransgrove et al (2025) also suggested that Hall waves, initiated following the superconducting transition in the core, may be sufficiently intense to fracture the crust, potentially leading to starquakes that induce rotational glitches and alterations in the radio pulse profile. Their simulations incorporated the time evolution of the elastic deformation of the lattice following previous work (Bransgrove et al, 2018), which enables them to monitor time-dependent shear stresses throughout the crust. Further investigations along these lines are worthwhile for future research.
@@ 3.5 Ambipolar diffusion in neutron star cores
The number of works concerning mechanisms operating in NS cores is sensibly smaller, and most contain far less detail than the studies of the crust. Ambipolar diffusion involves the coordinated movement of magnetic field lines and charged particles (protons and electrons) relative to neutrals (neutrons). There are several astrophysical scenarios, involving environments with partially ionized plasma like the Solar chromosphere, protostellar disks, and parts of the interstellar medium, where the ambipolar diffusion is regarded as a dominant mechanism. The seminal works by Goldreich and Reisenegger (1992) and Shalybkov and Urpin (1995), already proposed ambipolar diffusion as a viable mechanism for the dissipation of magnetic energy in regions of NSs where the charged particle fluid is chemically homogeneous. Owing to its cubic dependence on $B$, ambipolar diffusion could be the dominant process driving the evolution of magnetars during the first $10^{3}-10^{5}$ yr, although there is some controversy. In particular, we refer the reader interested in the role of chemical potential gradients, which is out of the scope of this review, to the literature, for instance the original arguments in Goldreich and Reisenegger (1992) or the discussion in Passamonti et al (2017). However, Gusakov et al (2017) questioned the validity of the approach followed by previous works in stratified matter, and obtained a different equation from the momentum equation (implicitly assuming magnetostatic equilibrium), in which the small deviations of the chemical potentials from their equilibrium values do not depend on temperature and are determined by the Lorentz force. With the same methodology, Ofengeim and Gusakov (2018) estimated the instantaneous particle velocities and other parameters of interest, determined by specifying the magnetic field configuration, and found that the evolution timescales could be shorter than expected.
A simple way to incorporate ambipolar diffusion is to introduce the “ambipolar velocity” $\vec{\mathrm{v}}_{a}$ and consider an advective term $\propto(-\vec{\mathrm{v}}_{a}\times\vec{B})$ in the electric field. As discussed in Goldreich and Reisenegger (1992); Passamonti et al (2017), the simplest case is realized in the regime where the system attains $\beta$-equilibrium faster than it evolves, and the ambipolar velocity is proportional to the Lorentz force
@@
${\vec{\mathrm{v}}_{a}}\times\vec{B}\propto[B^{2}\vec{j}-(\vec{j}\cdot\vec{B})\vec{B}]=B^{2}\vec{j}_{\perp}~,$ (28)
@@
Most previous studies of ambipolar diffusion have focused on estimating timescales, with only a few exceptions that include simulations. These simulations have primarily been limited to simplified one-dimensional models (Hoyos et al, 2008, 2010; Tsuruta et al, 2023). However, given the important considerations outlined above regarding the role of geometry, simplified one-dimensional results should be interpreted with caution, as they require confirmation through multidimensional studies. Two-dimensional (Castillo et al, 2017; Passamonti et al, 2017; Bransgrove et al, 2018; Skiathas and Gourgouliatos, 2024; Viganò et al, 2021), and the first three-dimensional simulations (Igoshev and Hollerbach, 2023), have become possible only very recently, although typically assuming constant coefficients and with some simplifying assumptions. Despite current numerical limitations, they already point to intriguing possibilities. A more consistent strategy based on a multifluid approach has recently gained attention, yielding promising results (Castillo et al, 2025; Moraga et al, 2025). Although still restricted to axisymmetric models, these initial findings support earlier estimates, indicating an enhancement in dissipation rates. At constant temperature, they recover the expected outcome: neutrons reach diffusive equilibrium, the Lorentz force is balanced by the chemical potential gradients of the charged particles, and the magnetic field configuration is governed by a non-linear Grad–Shafranov equation.
In addition, in a realistic scenario, there is a further complication that usually is omitted. The NS core cools very fast (less than a year) below the critical temperatures for neutron superfluidity and proton superconductivity, which has important implications, sometimes controversial. Goldreich and Reisenegger (1992) argued that ambipolar diffusion would still be a significant process, but Glampedakis et al (2011) studied in detail the ambipolar diffusion in superfluid and superconducting stars and concluded that its role on the magnetic field evolution would be negligible. Other works (Graber et al, 2015) also showed that ambipolar diffusion with superconducting protons is very slow. Moreover, Kantor and Gusakov (2018) argued that, in finite-temperature superfluid NS matter, magnetic field dissipates exclusively due to Ohmic losses and non-equilibrium beta-processes, limiting the effects to the case where muons are present.
@@ 3.6 Mathematical structure of the generalized induction equation
In order to understand the dynamical evolution of the system and to design a successful numerical algorithm, it is important to identify the mathematical character of the equations and the wave modes. The magnitude of $\omega_{B}^{e}\tau_{e}$ defines the transition from a purely parabolic equation ($\omega_{B}^{e}\tau_{e}\ll 1$) to a hyperbolic regime ($\omega_{B}^{e}\tau_{e}\gg 1$). The Hall term introduces two wave modes into the system. Huba (2003) has shown that, in a constant density medium, the only modes of the Hall–MHD equation are the whistler or helicon waves. They are transverse field perturbations propagating along the field lines. In presence of a charge density gradient, additional Hall drift waves appear. These are transverse modes that propagate in the $\vec{B}\times\vec{\nabla}n_{e}$ direction. We also note that the presence of charge density gradients results in a Burgers-like term (Vainshtein et al, 2000). Furthermore, even in the constant density case but without planar symmetry, the evolution of the toroidal component also contains a quadratic term that resembles the Burgers equation (Pons and Geppert, 2007) with a coefficient dependent on the distance to the axis. This term leads to the formation of discontinuous solutions (current sheets) that require proper treatment. It is fundamental for a numerical Hall–MHD code to reproduce these modes and features, which are easily testable.
@@
$\vec{B}=\vec{B}_{o}+\vec{\delta B}\,e^{i(\vec{k}\cdot\vec{x}-\omega t)}~,$ (30)
@@
and, following the standard procedure (see Viganò et al 2019, who did not include the CME term), the dispersion relation is given by
@@
In this form, we identify the ambipolar diffusion as a purely dissipative term, with a field-dependent diffusivity through $f_{a}\propto B_{0}^{2}$, but the CME term has opposite signs in each of the two modes, indicating that short-wavelength ($k<k_{5}/(1+f_{a})$) unstable modes are possible, if $k_{5}>0$ (see Eq. (24). Indeed, the mathematical form of the CME term, $\propto k_{5}\vec{B}$, is the same as the so-called $\alpha$-term in dynamo theory, although the CME pre-coefficient has a different physical origin and can have either sign.
@@ 4 Magnetic field evolution in the interior of neutron stars: numerical methods
Within spectral methods, another important choice is whether to apply a spectral decomposition in the radial direction (typically using Chebyshev polynomials) or to adopt a hybrid approach, combining spectral methods in angular directions with finite-difference schemes radially, which can more effectively resolve the steep density gradients in the NS crust. Examples of a fully spectral approach are the first 2D simulations of the evolution of the crustal magnetic field assumed a constant density shell (Hollerbach and Rüdiger, 2002) and were later extended to include density gradients (Hollerbach and Rüdiger, 2004). They used an adapted version of the spherical harmonic code described in Hollerbach (2000), including modes up to $l=100$, and 25 Chebyshev polynomials in the radial direction, but they were restricted to $\omega_{B}\tau_{e}<200$ by numerical issues.
@@
Pons and Geppert (2007); Pons et al (2009b) used a hybrid code (spectral in angles but finite-differences in the radial direction) to perform 2D simulations in realistic profiles of NSs over relevant timescales (typically, Myr). This approach allowed us to reach higher values of the magnetization parameter ($\omega_{B}\tau_{e}\approx 10^{3}$), and to study the Hall instability (Pons and Geppert, 2010). Similarly, a number of recent 3D simulations (Gourgouliatos and Hollerbach, 2018; Gourgouliatos and Pons, 2019; De Grandis et al, 2020, 2022; Igoshev et al, 2021, 2023), use a modified version of the PARODY code (Dormy et al, 1998; Aubert et al, 2008). An earlier version of this code, excluding thermomagnetic coupling, was already used in Wood and Hollerbach (2015) in the context of NSs. The code also employs a pseudo-spectral approach, with a radial grid and spherical harmonic expansions for the angular components. Time-stepping uses the Crank-Nicholson method for ohmic diffusion, backward Euler for the isotropic thermal diffusion, and Adams-Bashforth for all additional terms.
@@
Wood and Hollerbach (2015); Gourgouliatos et al (2016) were limited to magnetization parameters of the order of $\simeq 100$. The main problem arises from the presence of non-linear Burgers-like terms, which naturally lead to discontinuities (Vainshtein et al, 2000; Viganò et al, 2012), which are notoriously problematic for spectral codes. For this reason, subsequent works aiming at extending the simulations to more general cases have been gradually shifting towards the use of finite-difference schemes. An example of a finite difference approach, while still employing poloidal-toroidal decomposition, is seen in the axially symmetric simulations Gourgouliatos and Cumming (2014b, a, 2015). A more elaborated scheme using finite volume methods applied directly to the induction equation in terms of field components was presented in Viganò et al (2012) and used for different 2D applications Viganò and Pons (2012); Viganò et al (2013); Dehman et al (2020). They utilized Stokes’ theorem and the conservative form of the equations to apply techniques from high-resolution shock-capturing schemes, enabling handling of higher magnetization parameters. Extending these methods to 3D with a new code (MATINS) has only recently become feasible (Dehman et al, 2023c, b; Ascenzi et al, 2024; Dehman and Pons, 2025).
@@
It is worth noting that, despite differences in methods and coordinate choices, all existing codes handle the radial direction differently from the angular directions due to the stronger gradients along the radial axis. Alternatively, Cartesian grids can be used, but they present two main challenges: first, resolution cannot be enhanced solely in the radial direction, leading to a significant increase in computational cost compared to spherical coordinate-based codes (scaling as $\propto N^{3}$ instead of $N$, where $N$ is the number of radial points or dual basis components in spectral methods). Second, Cartesian discretization introduces numerical noise and spurious modes due to imperfect mapping of spherical boundaries. While these difficulties have recently been addressed in various star-in-a-box simulations in other contexts, the only existing preliminary application to neutron star evolution is presented in Viganò et al (2019), to which we refer for further details.
@@ 4.1 Spectral methods with the toroidal-poloidal decomposition
$\Phi=\frac{1}{r}\sum_{\ell,m}\Phi_{\ell m}(r,t)Y_{\ell m}(\theta,\varphi)~,$
$\Psi=\frac{1}{r}\sum_{\ell,m}\Psi_{\ell m}(r,t)Y_{\ell m}(\theta,\varphi)~,$ (36)
@@
Assuming a radial dependent diffusivity, $\eta=\eta(r)$, it can be shown that the Ohmic term for each multipole effectively decouples, and the set of coupled evolution equations for the radial parts ($\Phi_{\ell m}$ and $\Psi_{\ell m}$) can be readily obtained. Omitting relativistic factors (see Pons et al (2009a) for the relativistic expressions in the Schwarzschild metric) we have:
@@
Terms proportional to $k_{5}$ encode the chiral magnetic effect (App. A of Dehman and Pons 2025). For the quasi-analytical illustration we take $k_{5}$ constant; in general $k_{5}=k_{5}(\vec{r},t)$ is a spatially and temporally varying pseudo-scalar field. We use $D_{\ell m}$ and $C_{\ell m}$ as a shorthand for the nonlinear Hall terms (the full expressions can also be found in Geppert and Wiebicke 1991). These include sums over running indices and coupling constants related to Clebsch–Gordan coefficients (the sum rules to combine angular momentum operators are used to determine which multipoles are coupled to each other). All these coefficients can be evaluated once at the beginning of the evolution and stored in a memory-saving form since only specific combinations of indices are non-zero.

### v2

@@ 3.3 The chiral magnetic effect
Using a plane-parallel model, Lander and Gourgouliatos (2019) investigated the features of such plastic flow under the assumption of Stokes flow, where a viscous term balances magnetic and elastic stresses. They study the crustal response under Ohmic and Hall evolution and find that there can be significant plastic-like motions in the external layers of the star. Similar arguments have also been proposed to account for the deposition of heat by the visco-plastic flow and the propagation of thermo-plastic waves (Beloborodov and Levin, 2014). In a later work, Gourgouliatos and Lander (2021) conducted global axisymmetric simulations to investigate various failure mechanisms impacting a specific crustal region. They find that flow does not merely dampen the Hall effect, even when plastic viscosity is low; instead, it drives complex evolutionary patterns, sometimes amplifying the Hall effect’s influence. They concluded that plastic flow influences both the characteristics of magnetar bursts and their spin-down behavior.
However, as pointed out in recent works (Kojima, 2024; Bransgrove et al., 2025), in magneto-elastic equilibrium, it is the divergence of stress (i.e., the force) that is balanced, not the stress itself. Kojima’s methodology for calculating the elastic tensor is needed to ensure the model accuracy. In numerical simulations of magnetic field evolution, $\sigma_{ij}$ must be tracked by solving additional differential equations, at the cost of a significant increase in computation time compared to previous simple algebraic relations. Bransgrove et al. (2025) indicate that, in configurations close to magneto-elastic equilibrium, the elastic stress is typically several orders of magnitude smaller than the local Maxwell stress. Consequently, criteria based only on stress balance may significantly overestimate the frequency of crustal failures, and previous conclusions should be approached with caution. Nevertheless, the results in Bransgrove et al. (2025) also suggested that Hall waves, initiated following the superconducting transition in the core, may be sufficiently intense to fracture the crust, potentially leading to starquakes that induce rotational glitches and alterations in the radio pulse profile. Their simulations incorporated the time evolution of the elastic deformation of the lattice following previous work (Bransgrove et al., 2018), which enables them to monitor time-dependent shear stresses throughout the crust. Further investigations along these lines are worthwhile for future research.
@@ 3.5 Ambipolar diffusion in neutron star cores
The number of works concerning mechanisms operating in NS cores is sensibly smaller, and most contain far less detail than the studies of the crust. Ambipolar diffusion involves the coordinated movement of magnetic field lines and charged particles (protons and electrons) relative to neutrals (neutrons). There are several astrophysical scenarios, involving environments with partially ionized plasma like the Solar chromosphere, protostellar disks, and parts of the interstellar medium, where the ambipolar diffusion is regarded as a dominant mechanism. The seminal works by Goldreich and Reisenegger (1992) and Shalybkov and Urpin (1995), already proposed ambipolar diffusion as a viable mechanism for the dissipation of magnetic energy in regions of NSs where the charged particle fluid is chemically homogeneous. Owing to its cubic dependence on $B$, ambipolar diffusion could be the dominant process driving the evolution of magnetars during the first $10^{3}-10^{5}$ yr, although there is some controversy. In particular, we refer the reader interested in the role of chemical potential gradients, which is out of the scope of this review, to the literature, for instance the original arguments in Goldreich and Reisenegger (1992) or the discussion in Passamonti et al. (2017). However, Gusakov et al. (2017) questioned the validity of the approach followed by previous works in stratified matter, and obtained a different equation from the momentum equation (implicitly assuming magnetostatic equilibrium), in which the small deviations of the chemical potentials from their equilibrium values do not depend on temperature and are determined by the Lorentz force. With the same methodology, Ofengeim and Gusakov (2018) estimated the instantaneous particle velocities and other parameters of interest, determined by specifying the magnetic field configuration, and found that the evolution timescales could be shorter than expected.
A simple way to incorporate ambipolar diffusion is to introduce the “ambipolar velocity” $\vec{\mathrm{v}}_{a}$ and consider an advective term $\propto(-\vec{\mathrm{v}}_{a}\times\vec{B})$ in the electric field. As discussed in Goldreich and Reisenegger (1992); Passamonti et al. (2017), the simplest case is realized in the regime where the system attains $\beta$-equilibrium faster than it evolves, and the ambipolar velocity is proportional to the Lorentz force
@@
${\vec{\mathrm{v}}_{a}}\times\vec{B}\propto[B^{2}\vec{j}-(\vec{j}\cdot\vec{B})\vec{B}]=B^{2}\vec{j}_{\perp}\,$, (28)
@@
Most previous studies of ambipolar diffusion have focused on estimating timescales, with only a few exceptions that include simulations. These simulations have primarily been limited to simplified one-dimensional models (Hoyos et al., 2008, 2010; Tsuruta et al., 2023). However, given the important considerations outlined above regarding the role of geometry, simplified one-dimensional results should be interpreted with caution, as they require confirmation through multidimensional studies. Two-dimensional (Castillo et al., 2017; Passamonti et al., 2017; Bransgrove et al., 2018; Skiathas and Gourgouliatos, 2024; Viganò et al., 2021), and the first three-dimensional simulations (Igoshev and Hollerbach, 2023), have become possible only very recently, although typically assuming constant coefficients and with some simplifying assumptions. Despite current numerical limitations, they already point to intriguing possibilities. A more consistent strategy based on a multifluid approach has recently gained attention, yielding promising results (Castillo et al., 2025; Moraga et al., 2025). Although still restricted to axisymmetric models, these initial findings support earlier estimates, indicating an enhancement in dissipation rates. At constant temperature, they recover the expected outcome: neutrons reach diffusive equilibrium, the Lorentz force is balanced by the chemical potential gradients of the charged particles, and the magnetic field configuration is governed by a non-linear Grad–Shafranov equation.
In addition, in a realistic scenario, there is a further complication that usually is omitted. The NS core cools very fast (less than a year) below the critical temperatures for neutron superfluidity and proton superconductivity, which has important implications, sometimes controversial. Goldreich and Reisenegger (1992) argued that ambipolar diffusion would still be a significant process, but Glampedakis et al. (2011) studied in detail the ambipolar diffusion in superfluid and superconducting stars and concluded that its role on the magnetic field evolution would be negligible. Other works (Graber et al., 2015) also showed that ambipolar diffusion with superconducting protons is very slow. Moreover, Kantor and Gusakov (2018) argued that, in finite-temperature superfluid NS matter, magnetic field dissipates exclusively due to Ohmic losses and non-equilibrium beta-processes, limiting the effects to the case where muons are present.
@@ 3.6 Mathematical structure of the generalized induction equation
In order to understand the dynamical evolution of the system and to design a successful numerical algorithm, it is important to identify the mathematical character of the equations and the wave modes. The magnitude of $\omega_{B}^{e}\tau_{e}$ defines the transition from a purely parabolic equation ($\omega_{B}^{e}\tau_{e}\ll 1$) to a hyperbolic regime ($\omega_{B}^{e}\tau_{e}\gg 1$). The Hall term introduces two wave modes into the system. Huba (2003) has shown that, in a constant density medium, the only modes of the Hall–MHD equation are the whistler or helicon waves. They are transverse field perturbations propagating along the field lines. In presence of a charge density gradient, additional Hall drift waves appear. These are transverse modes that propagate in the $\vec{B}\times\vec{\nabla}n_{e}$ direction. We also note that the presence of charge density gradients results in a Burgers-like term (Vainshtein et al., 2000). Furthermore, even in the constant density case but without planar symmetry, the evolution of the toroidal component also contains a quadratic term that resembles the Burgers equation (Pons and Geppert, 2007) with a coefficient dependent on the distance to the axis. This term leads to the formation of discontinuous solutions (current sheets) that require proper treatment. It is fundamental for a numerical Hall–MHD code to reproduce these modes and features, which are easily testable.
@@
$\vec{B}=\vec{B}_{o}+\vec{\delta B}\,e^{i(\vec{k}\cdot\vec{x}-\omega t)}\,$, (30)
@@
and, following the standard procedure (see Viganò et al. 2019, who did not include the CME term), the dispersion relation is given by
@@
In this form, we identify the ambipolar diffusion as a purely dissipative term, with a field-dependent diffusivity through $f_{a}\propto B_{0}^{2}$, but the CME term has opposite signs in each of the two modes, indicating that short-wavelength ($k<k_{5}/(1+f_{a})$) unstable modes are possible, if $k_{5}>0$ (see Eq. 24). Indeed, the mathematical form of the CME term, $\propto k_{5}\vec{B}$, is the same as the so-called $\alpha$-term in dynamo theory, although the CME pre-coefficient has a different physical origin and can have either sign.
@@ 4 Magnetic field evolution in the interior of neutron stars: numerical methods
Within spectral methods, another important choice is whether to apply a spectral decomposition in the radial direction (typically using Chebyshev polynomials) or to adopt a hybrid approach, combining spectral methods in angular directions with finite-difference schemes radially, which can more effectively resolve the steep density gradients in the NS crust. Examples of a fully spectral approach are the first 2D simulations of the evolution of the crustal magnetic field assumed a constant density shell (Hollerbach and Rüdiger, 2002) and were later extended to include density gradients (Hollerbach and Rüdiger, 2004). They used an adapted version of the spherical harmonic code described in Hollerbach (2000), including modes up to $\ell=100$, and 25 Chebyshev polynomials in the radial direction, but they were restricted to $\omega_{B}\tau_{e}<200$ by numerical issues.
@@
Pons and Geppert (2007); Pons et al. (2009) used a hybrid code (spectral in angles but finite-differences in the radial direction) to perform 2D simulations in realistic profiles of NSs over relevant timescales (typically, Myr). This approach allowed us to reach higher values of the magnetization parameter ($\omega_{B}\tau_{e}\approx 10^{3}$), and to study the Hall instability (Pons and Geppert, 2010). Similarly, a number of recent 3D simulations (Gourgouliatos and Hollerbach, 2018; Gourgouliatos and Pons, 2019; De Grandis et al., 2020, 2022; Igoshev et al., 2021, 2023), use a modified version of the PARODY code (Dormy et al., 1998; Aubert et al., 2008). An earlier version of this code, excluding thermomagnetic coupling, was already used in Wood and Hollerbach (2015) in the context of NSs. The code also employs a pseudo-spectral approach, with a radial grid and spherical harmonic expansions for the angular components. Time-stepping uses the Crank-Nicholson method for ohmic diffusion, backward Euler for the isotropic thermal diffusion, and Adams-Bashforth for all additional terms.
@@
Wood and Hollerbach (2015); Gourgouliatos et al. (2016) were limited to magnetization parameters of the order of $\simeq 100$. The main problem arises from the presence of non-linear Burgers-like terms, which naturally lead to discontinuities (Vainshtein et al., 2000; Viganò et al., 2012), which are notoriously problematic for spectral codes. For this reason, subsequent works aiming at extending the simulations to more general cases have been gradually shifting towards the use of finite-difference schemes. An example of a finite difference approach, while still employing poloidal-toroidal decomposition, is seen in the axially symmetric simulations (Gourgouliatos and Cumming, 2014b, a, 2015). A more elaborated scheme using finite volume methods applied directly to the induction equation in terms of field components was presented in Viganò et al. (2012) and used for different 2D applications (Viganò and Pons, 2012; Viganò et al., 2013; Dehman et al., 2020). They utilized Stokes’ theorem and the conservative form of the equations to apply techniques from high-resolution shock-capturing schemes, enabling handling of higher magnetization parameters. Extending these methods to 3D with a new code (MATINS) has only recently become feasible (Dehman et al., 2023c, b; Ascenzi et al., 2024; Dehman and Pons, 2025).
@@
It is worth noting that, despite differences in methods and coordinate choices, all existing codes handle the radial direction differently from the angular directions due to the stronger gradients along the radial axis. Alternatively, Cartesian grids can be used, but they present two main challenges: first, resolution cannot be enhanced solely in the radial direction, leading to a significant increase in computational cost compared to spherical coordinate-based codes (scaling as $\propto N^{3}$ instead of $N$, where $N$ is the number of radial points or dual basis components in spectral methods). Second, Cartesian discretization introduces numerical noise and spurious modes due to imperfect mapping of spherical boundaries. While these difficulties have recently been addressed in various star-in-a-box simulations in other contexts, the only existing preliminary application to neutron star evolution is presented in Viganò et al. (2019), to which we refer for further details.
@@ 4.1 Spectral methods with the toroidal-poloidal decomposition
$\Phi=\frac{1}{r}\sum_{\ell,m}\Phi_{\ell m}(r,t)Y_{\ell m}(\theta,\varphi)\,$,
$\Psi=\frac{1}{r}\sum_{\ell,m}\Psi_{\ell m}(r,t)Y_{\ell m}(\theta,\varphi)\,$, (36)
@@
Assuming a radial dependent diffusivity, $\eta=\eta(r)$, it can be shown that the Ohmic term for each multipole effectively decouples, and the set of coupled evolution equations for the radial parts ($\Phi_{\ell m}$ and $\Psi_{\ell m}$) can be readily obtained. Omitting relativistic factors (see Pons et al. 2009 for the relativistic expressions in the Schwarzschild metric) we have:
@@
Terms proportional to $k_{5}$ encode the chiral magnetic effect (Appendix A of Dehman and Pons 2025). For the quasi-analytical illustration we take $k_{5}$ constant; in general $k_{5}=k_{5}(\vec{r},t)$ is a spatially and temporally varying pseudo-scalar field. We use $D_{\ell m}$ and $C_{\ell m}$ as a shorthand for the nonlinear Hall terms (the full expressions can also be found in Geppert and Wiebicke 1991). These include sums over running indices and coupling constants related to Clebsch–Gordan coefficients (the sum rules to combine angular momentum operators are used to determine which multipoles are coupled to each other). All these coefficients can be evaluated once at the beginning of the evolution and stored in a memory-saving form since only specific combinations of indices are non-zero.